# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 5.2. Direct limit systems at $\kappa_{i}$
For any $i\geq 0$, $\Sigma_{\mathcal{M}_{\infty}^{i}}\upharpoonright(\mathcal{M}|\delta_{i+1})\in\mathcal{M}.$ Moreover, if $\kappa_{i}$ is fully strong in $\mathcal{M}$, then $\Sigma_{\mathcal{M}_{\infty}^{i}}\upharpoonright\mathcal{M}\in\mathcal{M}.$
@@
As we are no longer working in the result of a fully backgrounded construction to obtain $\mathcal{M}_{\infty}^{i}$ but in the result of an $\mathcal{M}$-construction, we sketch how to modify the proof of the successor case in the proof of [12, Theorem 6.5] to obtain Lemma 5.11. We show the “moveover” part of the lemma parallelly to the first statement restricted to $\mathcal{M}|\delta_{i+1}$ as most of the proof is the same. We assume that the reader is familiar with the proof of [12, Theorem 6.5] and, for notational simplicity, restrict to the case for normal iteration trees. Sargsyan’s argument uses universality in the sense of [12, Lemma 2.13] or [23, Lemma 11.1] to, for example, show that $\mathcal{Q}$-structures for iteration trees according to the iteration strategy in question can be obtained via fully backgrounded constructions relative to segments of that strategy. This sort of universality goes back to Mitchell and Schindler’s proof of universality for $K^{c}$ in [7]. To apply Sargsyan’s argument in our setting we, for example, need the following claim. Before we can state the claim, recall the following notation from [12] (see also Subsection 3.1): For a hod premouse $\mathcal{P}$ with finitely many Woodin cardinals $\delta_{0},\dots,\delta_{i}$, write, if $i>0$, $\mathcal{P}^{-}=\mathcal{P}(i-1)$. We will also use this notation for initial segments of hod premice and, for example, write $(\mathcal{M}_{\infty}^{i,-})^{-}=(\mathcal{M}_{\infty}^{i})^{-}$. Note that $(\mathcal{M}_{\infty}^{i})^{-}$ is strictly shorter than $\mathcal{M}_{\infty}^{i,-}=\mathcal{M}_{\infty}^{i}|\delta^{\mathcal{M}_{\infty}^{i}}$.
@@
Let $\mathcal{T}$ be an iteration tree on $\mathcal{M}_{\infty}^{i}$ in $\mathcal{M}|\delta_{i+1}$ of limit length and let $b=\Sigma_{\mathcal{M}_{\infty}^{i}}(\mathcal{T})$. Suppose $\mathcal{T}$ is not entirely based on $(\mathcal{M}_{\infty}^{i})^{-}$ and $\mathcal{Q}(b,\mathcal{T})$ exists. Let $\mathcal{S}$ be the node in $\mathcal{T}$ such that the fragment of $\mathcal{T}$ from $\mathcal{M}_{\infty}^{i}$-to-$\mathcal{S}$ is based on $(\mathcal{M}_{\infty}^{i})^{-}$ and the rest of $\mathcal{T}$ is above $\mathcal{S}^{-}$. Let $\mathcal{N}$ be the result of a fully backgrounded construction relative to $(\Sigma_{(\mathcal{M}_{\infty}^{i})^{-}})_{\mathcal{S}^{-}}$ above $\mathcal{M}(\mathcal{T})$ in $\mathcal{M}|\delta_{i+1}$. Then
@@
    the comparison of $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ in $\mathcal{M}|\delta_{i+1}$ is successful, and
@@
Note that (2) follows from (1) by universality of fully backgrounded constructions as in [12, Lemma 2.13] and [23, Lemma 11.1]. In particular, it suffices for universality of $\mathcal{N}$ against $\mathcal{Q}(b,\mathcal{T})$ that the comparison of $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ is successful in $\mathcal{M}|\delta_{i+1}$. So we are left with showing (1).
Suppose toward a contradiction that the comparison is not successful and let $\mathcal{U}_{0}$ and $\mathcal{U}_{1}$ be iteration trees on $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ witnessing this. Take a countable substructure of a sufficiently large initial segment of $\mathcal{M}$ reflecting this situation and write $\bar{\mathcal{Q}}(b,\mathcal{T}),\bar{\mathcal{N}},\bar{\mathcal{U}}_{0}$, and $\bar{\mathcal{U}}_{1}$ for the images of $\mathcal{Q}(b,\mathcal{T}),\mathcal{N},\mathcal{U}_{0}$, and $\mathcal{U}_{1}$ under the collapse. By the iterability proof for fully backgrounded constructions, the relevant iterates of $\bar{\mathcal{N}}$ can be realized into $\mathcal{M}$. Moreover, the $\bar{\mathcal{N}}$-side of the comparison provides $\mathcal{Q}$-structures for the $\bar{\mathcal{Q}}(b,\mathcal{T})$-side, cf., for example [12, Fact 3.6]. Therefore, the comparison of $\bar{\mathcal{Q}}(b,\mathcal{T})$ and $\bar{\mathcal{N}}$ is successful in $\mathcal{M}$. ∎
A similar claim can be shown for $\mathcal{T}\in\mathcal{M}$ for the “moreover” part of Lemma 5.11. Using the notation in Claim 1, the cases in the proof of [12, Theorem 6.5] when the tree $\mathcal{T}$ is entirely based on $(\mathcal{M}_{\infty}^{i})^{-}$ and when there is a fatal drop in the rest of $\mathcal{T}$ above $\mathcal{S}^{-}$ are dealt with similarly as the case where $\mathcal{Q}(b,\mathcal{T})$ exists. So we are left with showing that branches for maximal trees are also computed correctly (this corresponds to Clause $2(c)$ in the proof of [12, Theorem 6.5]).
We define the iteration strategy $\Lambda$ as in the proof of [12, Theorem 6.5]. This means that given $\mathcal{T}\in\operatorname{dom}(\Lambda)$ such that $\mathcal{T}$ is not covered by one of the cases discussed above, we say $\Lambda(\mathcal{T})$ is defined and let $\Lambda(\mathcal{T})=b$ iff there is an extender $E$ on the sequence of $\mathcal{M}$ with critical point $\kappa_{i}$ such that $\mathcal{T}\in\mathcal{M}|\operatorname{lh}(E)$ and there is an embedding
$\sigma\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{M}_{\infty}^{i})$
such that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. As in the proof of [12, Theorem 6.5] we have that $\mathcal{T}\in\operatorname{dom}(\Lambda)$ if and only if $\mathcal{T}\in\operatorname{dom}(\Sigma_{\mathcal{M}_{\infty}^{i}})$. We are left with showing the following claim.
@@
Let $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{M}_{\infty}^{i}})$ be as above. Then
@@
    $\Lambda(\mathcal{T})$ is defined for $\mathcal{T}\in\mathcal{M}|\delta_{i+1}$, and
@@
    if it is defined, $\Lambda(\mathcal{T})=\Sigma_{\mathcal{M}_{\infty}^{i}}(\mathcal{T})$.
Moreover, if we inductively suppose that $\kappa_{i}$ is a fully strong cardinal in $\mathcal{M}$, then $\Lambda(\mathcal{T})$ is defined for all $\mathcal{T}\in\mathcal{M}$.
@@
For (2), let $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{M}_{\infty}^{i}})$ be such that $\Lambda(\mathcal{T})$ is defined. So let $\Lambda(\mathcal{T})=b$ and let $E$ be an extender on the sequence of $\mathcal{M}$ with critical point $\kappa_{i}$ and let $\sigma\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{M}_{\infty}^{i})$ be an embedding such that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. In case the extender $E$ is added to $\mathcal{M}$ by the fully backgrounded part of the construction (which will be the case if $\operatorname{lh}(E)<\delta_{i+1}$) the argument in the proof of [12, Theorem 6.5] applies. So assume that is not the case and $E$ is generically countably complete.
@@
Let $\hat{\mathcal{W}}$ be a countable elementary substructure of a sufficiently large initial segment $\mathcal{W}|\Omega$ of $\mathcal{W}$. Let $\tau\colon\hat{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$ and suppose that all relevant objects are in the range of $\tau$. For notational simplicity write $\mathcal{M}$ for the result of the $\mathcal{M}$-construction in $\mathcal{W}|\Omega$ and let $\hat{\mathcal{M}}$ be the transitive collapse of $\mathcal{M}$, i.e., the result of the $\mathcal{M}$-construction in $\hat{\mathcal{W}}$. Say $\tau(\hat{\kappa}_{i},\hat{E},\hat{\mathcal{M}}_{\infty}^{i},\hat{\Sigma}^{i,\infty})=(\kappa_{i},E,\mathcal{M}_{\infty}^{i},\Sigma_{\mathcal{M}_{\infty}^{i}})$.
@@
When comparing ^$\hat{\mathcal{M}}_{\infty}^{i}$ via $\hat{\Sigma}^{i,\infty}$ and $\mathcal{M}_{\infty}^{i}$ via $\Sigma_{\mathcal{M}_{\infty}^{i}}$ inside $\mathcal{W}$, ^$\hat{\mathcal{M}}_{\infty}^{i}$ iterates to $\mathcal{M}_{\infty}^{i}$. Moreover,
$\tau\upharpoonright\hat{\mathcal{M}}_{\infty}^{i}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{M}}_{\infty}^{i},\tau(\hat{\mathcal{M}}_{\infty}^{i})}$
@@
as there are term relations that are strongly guiding $\hat{\Sigma}^{i,\infty}$. We have that $E=\tau(\hat{E})$ is generically countably complete, so by Fact 4.2 there is a realization map $r\colon\operatorname{Ult}(\hat{\mathcal{M}},\hat{E})\rightarrow\mathcal{M}$ such that $\tau\upharpoonright\hat{\mathcal{M}}=r\circ\pi_{\hat{E}}$ where $\pi_{\hat{E}}\colon\hat{\mathcal{M}}\rightarrow\operatorname{Ult}(\hat{\mathcal{M}},\hat{E})$ denotes the ultrapower embedding. Therefore, $\pi_{\hat{E}}$ respects $\hat{\Sigma}^{i,\infty}$ and $\pi_{\hat{E}}\upharpoonright\hat{\mathcal{M}}_{\infty}^{i}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{M}}_{\infty}^{i},\pi_{\hat{E}}(\hat{\mathcal{M}}_{\infty}^{i})}.$ By elementarity of $\tau$,
$\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\pi^{\Sigma_{\mathcal{M}_{\infty}^{i}}}_{\mathcal{M}_{\infty}^{i},\pi_{E}(\mathcal{M}_{\infty}^{i})}$
@@
and hence $\pi^{\Sigma_{\mathcal{M}_{\infty}^{i}}}_{\mathcal{M}_{\infty}^{i},\pi_{E}(\mathcal{M}_{\infty}^{i})}=\sigma\circ\pi_{b}^{\mathcal{T}}$. As $\Sigma_{\mathcal{M}_{\infty}^{i}}$ has branch condensation, $b=\Sigma_{\mathcal{M}_{\infty}^{i}}(\mathcal{T})$, as desired.
@@
For (1), given $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{M}_{\infty}^{i}})$ as above and $b=\Sigma_{\mathcal{M}_{\infty}^{i}}(\mathcal{T})$ we aim to find $E$ and $\sigma$ such that $\pi_{E}\upharpoonright\mathcal{M}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. Let $r<\omega$ with $r\geq i+1$ be least such that $\mathcal{T}\in\mathcal{M}|\delta_{r}$. We start with arguing that $\mathcal{M}_{b}^{\mathcal{T}}$ is in fact an element of $\mathcal{M}$. Let $\mathcal{N}$ be the result of a fully backgrounded construction relative to $(\Sigma_{\mathcal{M}_{b}^{\mathcal{T}}})^{-}$ above $\mathcal{M}(\mathcal{T})$ in $\mathcal{M}|\delta_{r}$ using extenders with critical points above $\kappa_{r-1}$. Recall that inductively $(\Sigma_{\mathcal{M}_{b}^{\mathcal{T}}})^{-}\upharpoonright\mathcal{M}\in\mathcal{M}$. As the extenders backgrounding the construction of $\mathcal{N}$ in $\mathcal{M}|\delta_{r}$ have critical points above $\kappa_{r-1}$ they are in fact fully backgrounded in $\mathcal{W}$. Therefore, when comparing $\mathcal{M}_{b}^{\mathcal{T}}$ against the construction of $\mathcal{N}$ in $\mathcal{W}$, by universality and stationarity of fully backgrounded constructions [12, Lemmas 2.11 and 2.13], $\mathcal{M}_{b}^{\mathcal{T}}$ iterates to a level of the construction of $\mathcal{N}$. By fullness preservation of the iteration from $\mathcal{P}^{i}$ via $\mathcal{M}_{\infty}^{i}$ to $\mathcal{M}_{b}^{\mathcal{T}}$, the iteration from $\mathcal{M}_{b}^{\mathcal{T}}$ to the construction of $\mathcal{N}$ would have to drop if it is non-trivial since $\mathcal{M}_{b}^{\mathcal{T}}$ and $\mathcal{N}$ agree up to $\delta(\mathcal{T})$. Hence, $\mathcal{M}_{b}^{\mathcal{T}}\unlhd\mathcal{N}$ and, in particular, $\mathcal{M}_{b}^{\mathcal{T}}\in\mathcal{M}$.
@@
In case $\mathcal{T}\in\mathcal{M}|\delta_{i+1}$ recall that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}$ and pick $E$ with critical point $\kappa_{i}$ and $\mathcal{T}\in\mathcal{M}|\operatorname{lh}(E)$ on the sequence of $\mathcal{M}$ added by the fully backgrounded part of the $\mathcal{M}$-construction. In this case, $r=i+1$. For the “moreover” part of the claim, we suppose inductively that $\kappa_{i}$ is fully strong in $\mathcal{M}$, so we can fix an extender $E$ on the sequence of $\mathcal{M}$ with $\operatorname{lh}(E)>\delta_{r}$. In both cases, we pick $E$ such that, in addition to the requirements above, $\mathcal{M}$ and $\pi_{E}(\mathcal{M})$ agree on the result $\mathcal{N}$ of the hybrid fully backgrounded construction up to $\mathcal{M}_{b}^{\mathcal{T}}$. As most of what follows will be the same for both cases, we will deal with them parallelly.

### v3

@@ 5.2. Direct limit systems at $\kappa_{i}$
Let $k\geq 0$ and suppose $(\mathsf{IH})_{k}$. Then, for any $i\leq k$, $\Sigma_{\mathcal{P}_{\infty}^{i}}\upharpoonright\mathcal{M}^{(k)}|\delta_{k+1}$ is amenable to $\mathcal{M}^{(k)}|\delta_{k+1}$.
@@
As we are no longer working in the result of a fully backgrounded construction to obtain $\mathcal{P}_{\infty}^{i}$ but in the result of an $\mathcal{M}$-construction, we sketch how to modify the proof of the successor case in the proof of [13, Theorem 6.5] to obtain Lemma 5.12. We assume that the reader is familiar with the proof of [13, Theorem 6.5] and, for notational simplicity, restrict to the case for normal iteration trees. Sargsyan’s argument uses universality in the sense of [13, Lemma 2.13] or [25, Lemma 11.1] to, for example, show that $\mathcal{Q}$-structures for iteration trees according to the iteration strategy in question can be obtained via fully backgrounded constructions relative to segments of that strategy. This sort of universality goes back to Mitchell and Schindler’s proof of universality for $K^{c}$ in [7]. To apply Sargsyan’s argument in our setting we, for example, need the following claim. Before we can state the claim, recall the following notation from [13] (see also Subsection 3.1): For a hod premouse $\mathcal{P}$ with finitely many Woodin cardinals $\delta_{0},\dots,\delta_{n}$, write, if $n>0$, $\mathcal{P}^{-}=\mathcal{P}(n-1)$. We will also use this notation for initial segments of hod premice and, for example, write $(\mathcal{P}_{\infty}^{i,-})^{-}=(\mathcal{P}_{\infty}^{i})^{-}$. Note that $(\mathcal{P}_{\infty}^{i})^{-}$ is strictly shorter than $\mathcal{P}_{\infty}^{i,-}=\mathcal{P}_{\infty}^{i}|\delta^{\mathcal{P}_{\infty}^{i}}$. Recall that $(\mathsf{IH})_{i}$ implies that $\mathcal{M}^{(i)}|\delta_{i+1}$ is well-defined.
@@
Let $\mathcal{T}$ be an iteration tree on $\mathcal{P}_{\infty}^{i}$ in $\mathcal{M}^{(i)}|\delta_{i+1}$ of limit length and let $b=\Sigma_{\mathcal{P}_{\infty}^{i}}(\mathcal{T})$. Suppose $\mathcal{T}$ is not entirely based on $(\mathcal{P}_{\infty}^{i})^{-}$ and $\mathcal{Q}(b,\mathcal{T})$ exists. Let $\mathcal{S}$ be the node in $\mathcal{T}$ such that the fragment of $\mathcal{T}$ from $\mathcal{P}_{\infty}^{i}$-to-$\mathcal{S}$ is based on $(\mathcal{P}_{\infty}^{i})^{-}$ and the rest of $\mathcal{T}$ is above $\mathcal{S}^{-}$. Let $\mathcal{N}$ be the result of a fully backgrounded construction relative to $(\Sigma_{(\mathcal{P}_{\infty}^{i})^{-}})_{\mathcal{S}^{-}}$ above $\mathcal{M}(\mathcal{T})$ in $\mathcal{M}^{(i)}|\delta_{i+1}$. Then
@@
    the comparison of $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ in $\mathcal{M}^{(i)}|\delta_{i+1}$ is successful, and
@@
Note that (2) follows from (1) by universality of fully backgrounded constructions as in [13, Lemma 2.13] and [25, Lemma 11.1]. In particular, it suffices for universality of $\mathcal{N}$ against $\mathcal{Q}(b,\mathcal{T})$ that the comparison of $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ is successful in $\mathcal{M}^{(i)}|\delta_{i+1}$. So we are left with showing (1).
Suppose toward a contradiction that the comparison is not successful and let $\mathcal{U}_{0}$ and $\mathcal{U}_{1}$ be iteration trees on $\mathcal{Q}(b,\mathcal{T})$ and $\mathcal{N}$ witnessing this. Take a countable substructure of a sufficiently large initial segment of $\mathcal{M}$ reflecting this situation and write $\bar{\mathcal{Q}}(b,\mathcal{T}),\bar{\mathcal{N}},\bar{\mathcal{U}}_{0}$, and $\bar{\mathcal{U}}_{1}$ for the images of $\mathcal{Q}(b,\mathcal{T}),\mathcal{N},\mathcal{U}_{0}$, and $\mathcal{U}_{1}$ under the collapse. By the iterability proof for fully backgrounded constructions, the relevant iterates of $\bar{\mathcal{N}}$ can be realized into $\mathcal{M}$. Moreover, the $\bar{\mathcal{N}}$-side of the comparison provides $\mathcal{Q}$-structures for the $\bar{\mathcal{Q}}(b,\mathcal{T})$-side, cf., for example [13, Fact 3.6]. Therefore, the comparison of $\bar{\mathcal{Q}}(b,\mathcal{T})$ and $\bar{\mathcal{N}}$ is successful in $\mathcal{M}$. ∎
A similar claim can be shown for $\mathcal{T}\in\mathcal{M}^{(k)}|\delta_{k+1}$ for the case $i<k$ in the statement of Lemma 5.12, using that by $(\mathsf{IH})_{k}$, $\mathcal{M}^{(k)}|\delta_{k+1}$ is a well-defined model. Using the notation in Claim 1, the cases in the proof of [13, Theorem 6.5] when the tree $\mathcal{T}$ is entirely based on $(\mathcal{P}_{\infty}^{i})^{-}$ and when there is a fatal drop in the rest of $\mathcal{T}$ above $\mathcal{S}^{-}$ are dealt with similarly as the case where $\mathcal{Q}(b,\mathcal{T})$ exists. So we are left with showing that branches for maximal trees are also computed correctly (this corresponds to Clause $2(c)$ in the proof of [13, Theorem 6.5]).
We define the iteration strategy $\Lambda$ as in the proof of [13, Theorem 6.5]. This means that given $\mathcal{T}\in\operatorname{dom}(\Lambda)$ such that $\mathcal{T}$ is not covered by one of the cases discussed above, we say $\Lambda(\mathcal{T})$ is defined and let $\Lambda(\mathcal{T})=b$ iff there is an extender $E$ on the sequence of $\mathcal{M}^{(k)}|\delta_{k+1}$ with critical point $\kappa_{i}$ such that $\mathcal{T}\in\mathcal{M}^{(k)}|\operatorname{lh}(E)$ and there is an embedding
$\sigma\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{P}_{\infty}^{i})$
such that $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. As in the proof of [13, Theorem 6.5] we have that $\mathcal{T}\in\operatorname{dom}(\Lambda)$ if and only if $\mathcal{T}\in\operatorname{dom}(\Sigma_{\mathcal{P}_{\infty}^{i}})$. We are left with showing the following claim.
@@
Let $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{P}_{\infty}^{i}})$ be as above. Then
@@
    $\Lambda(\mathcal{T})$ is defined for $\mathcal{T}\in\mathcal{M}^{(k)}|\delta_{k+1}$, and
@@
    if it is defined, $\Lambda(\mathcal{T})=\Sigma_{\mathcal{P}_{\infty}^{i}}(\mathcal{T})$.
@@
For (2), let $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{P}_{\infty}^{i}})$ be such that $\Lambda(\mathcal{T})$ is defined. So let $\Lambda(\mathcal{T})=b$ and let $E$ be an extender on the sequence of $\mathcal{M}^{(k)}|\delta_{k+1}$ with critical point $\kappa_{i}$ and let $\sigma\colon\mathcal{M}_{b}^{\mathcal{T}}\rightarrow\pi_{E}(\mathcal{P}_{\infty}^{i})$ be an embedding such that $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. In case the extender $E$ is added to $\mathcal{M}^{(k)}|\delta_{k+1}$ by the fully backgrounded part of the construction (which will be the case if $\operatorname{lh}(E)<\delta_{i+1}$) the argument in the proof of [13, Theorem 6.5] applies. So assume that is not the case and $E$ is generically countably complete.
@@
Let $\hat{\mathcal{W}}$ be a countable elementary substructure of a sufficiently large initial segment $\mathcal{W}|\Omega$ of $\mathcal{W}$. Let $\tau\colon\hat{\mathcal{W}}\rightarrow\mathcal{W}|\Omega$ and suppose that all relevant objects are in the range of $\tau$. For notational simplicity write $\mathcal{M}^{(k)}|\delta_{k+1}$ for the result of the $\mathcal{M}$-construction up to $\delta_{k+1}$ where Clause (2c) is restricted to extenders with critical point in $\{\kappa_{j}\mid j<k\}$ in $\mathcal{W}|\Omega$ and let $\hat{\mathcal{M}}$ be the transitive collapse of $\mathcal{M}^{(k)}|\delta_{k+1}$, i.e., the result of the same type of $\mathcal{M}$-construction in $\hat{\mathcal{W}}$. Say $\tau(\hat{\kappa}_{i},\hat{E},\hat{\mathcal{P}}_{\infty}^{i},\hat{\Sigma}^{i,\infty})=(\kappa_{i},E,\mathcal{P}_{\infty}^{i},\Sigma_{\mathcal{P}_{\infty}^{i}})$.
@@
When comparing $\hat{\mathcal{P}}_{\infty}^{i}$ via $\hat{\Sigma}^{i,\infty}$ and $\mathcal{P}_{\infty}^{i}$ via $\Sigma_{\mathcal{P}_{\infty}^{i}}$ inside $\mathcal{W}$, $\hat{\mathcal{P}}_{\infty}^{i}$ iterates to $\mathcal{P}_{\infty}^{i}$. Moreover,
$\tau\upharpoonright\hat{\mathcal{P}}_{\infty}^{i}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{P}}_{\infty}^{i},\tau(\hat{\mathcal{P}}_{\infty}^{i})}$
@@
as there are term relations that are strongly guiding $\hat{\Sigma}^{i,\infty}$. We have that $E=\tau(\hat{E})$ is generically countably complete, so by Fact 4.2 there is a realization map $r\colon\operatorname{Ult}(\hat{\mathcal{M}},\hat{E})\rightarrow\mathcal{M}$ such that $\tau\upharpoonright\hat{\mathcal{M}}=r\circ\pi_{\hat{E}}$ where $\pi_{\hat{E}}\colon\hat{\mathcal{M}}\rightarrow\operatorname{Ult}(\hat{\mathcal{M}},\hat{E})$ denotes the ultrapower embedding. Therefore, $\pi_{\hat{E}}$ respects $\hat{\Sigma}^{i,\infty}$ and $\pi_{\hat{E}}\upharpoonright\hat{\mathcal{P}}_{\infty}^{i}=\pi^{\hat{\Sigma}^{i,\infty}}_{\hat{\mathcal{P}}_{\infty}^{i},\pi_{\hat{E}}(\hat{\mathcal{P}}_{\infty}^{i})}.$ By elementarity of $\tau$,
$\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\pi^{\Sigma_{\mathcal{P}_{\infty}^{i}}}_{\mathcal{P}_{\infty}^{i},\pi_{E}(\mathcal{P}_{\infty}^{i})}$
@@
and hence $\pi^{\Sigma_{\mathcal{P}_{\infty}^{i}}}_{\mathcal{P}_{\infty}^{i},\pi_{E}(\mathcal{P}_{\infty}^{i})}=\sigma\circ\pi_{b}^{\mathcal{T}}$. As $\Sigma_{\mathcal{P}_{\infty}^{i}}$ has branch condensation, $b=\Sigma_{\mathcal{P}_{\infty}^{i}}(\mathcal{T})$, as desired.
@@
For (1), given $\mathcal{T}\in\operatorname{dom}(\Lambda)\cap\operatorname{dom}(\Sigma_{\mathcal{P}_{\infty}^{i}})$ as above and $b=\Sigma_{\mathcal{P}_{\infty}^{i}}(\mathcal{T})$ we aim to find $E$ and $\sigma$ such that $\pi_{E}\upharpoonright\mathcal{P}_{\infty}^{i}=\sigma\circ\pi_{b}^{\mathcal{T}}$. Let $r\leq k$ with $r\geq i+1$ be least such that $\mathcal{T}\in\mathcal{M}^{(r)}|\delta_{r}$. We start with arguing that $\mathcal{M}_{b}^{\mathcal{T}}$ is in fact an element of $\mathcal{M}^{(r)}$. Let $\mathcal{N}$ be the result of a fully backgrounded construction relative to $(\Sigma_{\mathcal{M}_{b}^{\mathcal{T}}})^{-}$ above $\mathcal{M}(\mathcal{T})$ in $\mathcal{M}^{(r)}|\delta_{r}$ using extenders with critical points above $\kappa_{r-1}$. Recall that inductively $(\Sigma_{\mathcal{M}_{b}^{\mathcal{T}}})^{-}\upharpoonright(\mathcal{M}^{(r)}|\delta_{r})$ is amenable to $\mathcal{M}^{(r)}|\delta_{r}$. As the extenders backgrounding the construction of $\mathcal{N}$ in $\mathcal{M}^{(r)}|\delta_{r}$ have critical points above $\kappa_{r-1}$ they are in fact fully backgrounded in $\mathcal{W}$. Therefore, when comparing $\mathcal{M}_{b}^{\mathcal{T}}$ against the construction of $\mathcal{N}$ in $\mathcal{W}$, by universality and stationarity of fully backgrounded constructions [13, Lemmas 2.11 and 2.13], $\mathcal{M}_{b}^{\mathcal{T}}$ iterates to a level of the construction of $\mathcal{N}$. By fullness preservation of the iteration from $\mathcal{P}^{i}$ via $\mathcal{P}_{\infty}^{i}$ to $\mathcal{M}_{b}^{\mathcal{T}}$, the iteration from $\mathcal{M}_{b}^{\mathcal{T}}$ to the construction of $\mathcal{N}$ would have to drop if it is non-trivial since $\mathcal{M}_{b}^{\mathcal{T}}$ and $\mathcal{N}$ agree up to $\delta(\mathcal{T})$. Hence, $\mathcal{M}_{b}^{\mathcal{T}}\unlhd\mathcal{N}$ and, in particular, $\mathcal{M}_{b}^{\mathcal{T}}\in\mathcal{M}^{(r)}|\delta_{r}$.
@@
In case $i=k$, i.e., $\mathcal{T}\in\mathcal{M}^{(i)}|\delta_{i+1}$, recall that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong in $\mathcal{M}^{(i)}$ and pick $E$ with critical point $\kappa_{i}$ and $\mathcal{T}\in\mathcal{M}^{(i)}|\operatorname{lh}(E)$ on the sequence of $\mathcal{M}^{(i)}$ added by the fully backgrounded part of the $\mathcal{M}$-construction. In this case, $r=i+1$. For the case $k>i$, we suppose inductively that $(\mathsf{IH})_{k}$ holds, so $\kappa_{i}$ is strong in $\mathcal{M}^{(k)}|\delta_{k+1}$. Therefore, we can fix an extender $E$ on the sequence of $\mathcal{M}^{(k)}|\delta_{k+1}$ with $\operatorname{lh}(E)>\delta_{r}$. In both cases, we pick $E$ such that, in addition to the requirements above, $\mathcal{M}^{(k)}|\delta_{k+1}$ and $\pi_{E}(\mathcal{M}^{(k)}|\delta_{k+1})$ agree on the result $\mathcal{N}$ of the hybrid fully backgrounded construction up to $\mathcal{M}_{b}^{\mathcal{T}}$. As most of what follows will be the same for both cases, $k=i$ and $k>i$, we will deal with them parallelly.